OpenStudy (anonymous):

If y = 5/2(e^(x/5) + e^(-x/5)), prove that y'' = y/25.

OpenStudy (anonymous):

I expanded the bracket: y = 5/2(e^(x/5)) + 5/2(e^(-x/5))

OpenStudy (anonymous):

Then I took the first derivative:

OpenStudy (anonymous):

y ' = 5x/10(e^(x/5)) - 5x/10(e^(-x/5))

OpenStudy (anonymous):

Now how should I go about with getting the second derivative? Should I factor out the 5x/10 and then use product rule?

OpenStudy (anonymous):

\[y'=5/2[e ^{x/5}(1/5)+e^{-x/5}(-1/5)\]

OpenStudy (anonymous):

\[y'=(5/2)(1/5)[e ^{x/5}-e ^{-x/5}]\]

OpenStudy (anonymous):

Does it make a difference if we were to expand first with 5/2, and then take the first derivaitve?

OpenStudy (anonymous):

not at all

OpenStudy (anonymous):

Because that is what I did and then I took the dy/dx of each of the two terms.

OpenStudy (anonymous):

but y did u multiply with x/5?

OpenStudy (anonymous):

instead it is 1/5 the coefficient of x/5

OpenStudy (anonymous):

No sorry I did the following: y = 5/2(e^{x/5}) + 5/2(e^{-x/5}) y '= 5/2(1/5)*(e^{x/5}) + 5/2(-1/5)(e^{-x/5})

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

So then: y = 1/2(e^{x/5} + e^{-x/5})

OpenStudy (anonymous):

y' u mean?

OpenStudy (anonymous):

yes sorry, y'

OpenStudy (anonymous):

y' = 1/2(e^{x/5} + e^{-x/5})

OpenStudy (anonymous):

n there must b negative between to terms

OpenStudy (anonymous):

coz 1/5 is the common factor

OpenStudy (anonymous):

y' = 1/2(e^{x/5} - e^{-x/5}), right?

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

So then for y'' I get..

OpenStudy (anonymous):

\[y''=1/2[e ^{^{x/5}}(1/5)-e ^{-x/5}(-1/5)]\]

OpenStudy (anonymous):

Yes I got that, then I expanded, and I got: y'' = 1/10(e^{x/5} + e^{-x/5})

OpenStudy (anonymous):

\[y''=1/10[e ^{x/5}+e ^{-x/5}]\]

OpenStudy (anonymous):

again no need to expand

OpenStudy (anonymous):

\[e ^{x/5}+ e ^{-x/5}=2y/5\]

OpenStudy (anonymous):

y''=y/25

OpenStudy (anonymous):

How did you get e^x/5 + e^-x/5 = 2y/5?

OpenStudy (anonymous):

OHHH

OpenStudy (anonymous):

I think I know

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

I got it thank you so much. So when you are taking derivative of a function that has a constant which is multipled by a bracket that can be differentiated, you just take the derivative of the terms inside the bracket and multiply by the constant in front of it?

OpenStudy (anonymous):

right:)

OpenStudy (anonymous):

Ok awesome. Thank you for your help.

OpenStudy (anonymous):

my pleasure